OpenStudy (anonymous):

Solve 2(3x – 4) = 3x + 7. i must show all work, so please help:)

OpenStudy (anonymous):

2(3x – 4) = 3x + 7 6x - 8 = 3x + 7 3x - 8 + 8 = 7 + 8 3x = 15 x = 5

OpenStudy (anonymous):

thanks so much:)

OpenStudy (anonymous):

no probs. just be sure to help the others too if you can.

OpenStudy (anonymous):

i'll try! people respond so quick though..

OpenStudy (anonymous):

Trying is more than enough

OpenStudy (anonymous):

exactly :)